# The Basic Reproduction Number for Petri Net Models: A Next-Generation Matrix Approach

### v2

@@ 3.2 Next-Generation Matrix for Petri Nets
Figure 1: This VAPN is used as a guide to help visualize some of the concepts described in the Next-Generation Matrix for Petri Nets formulation. The places of this VAPN can be labeled $(P_{1},P_{2},P_{3},P_{4},P_{5})$ or equivalently $(N_{1},I_{1},I_{2},I_{3},N_{2})$, where the second label gives more description on the infectious state of the place.
@@
First, think of places in Petri nets the same way that you might think about compartments in ODEs. Then we assume we have a Petri net model for a disease that fits the definition of VAPN or SPN laid out in Section 2. Assume that each transition can fire once per time step. Define the total amount infected places in the Petri net model as $T$. Thus, the infected places are defined as $I_{1},I_{2},...,I_{T}$. As a note, we define all places as $P$ and they are labeled $P_{1},...,P_{L}$, where $L$ is total number of places and $T\leq L$. We define $t_{N}$ as a transition with an input arc coming from a non-infected place and an output arc going to an infected place. With $w(t_{N_{j,...,k}},I_{i})$ being the sum of all arcs weights $j,...,k$ from arcs going from non-infected places going to an infected place $I_{i}$ for VAPN or the net sum of all rate equations $j,...,k$ for arcs from transitions with input arcs from non-infected places and output arcs going to infected places $I_{i}$ for SPN. Let $x=(x_{1},…,x_{n})^{t}$, with each $x_{i}\geq 0$, be the number of tokens in each place, being the Petri net equivalent of population in each compartment in an ODE. Then we define $\mathcal{F}_{i}(x)$ to be the rate at which tokens move into infected places. For the variable arc weight PN, this rate refers to the arc weights. This is represented in matrix form by
$\mathcal{F}_{1,...,T}(x)$ $=$ $\begin{bmatrix}w(t_{N_{j,...,k}},I_{1})&w(t_{N_{j,...,k}},I_{1})&\cdots&w(t_{N_{j,...,k}},I_{1})\\ w(t_{N_{j,...,k}},I_{2})&w(t_{N_{j,...,k}},I_{2})&\cdots&w(t_{N_{j,...,k}},I_{2})\\ \vdots&\vdots&\ddots&\vdots\\ w(t_{N_{j,...,k}},I_{T})&w(t_{N_{j,...,k}},I_{T})&\cdots&w(t_{N_{j,...,k}},I_{T}))\end{bmatrix}.$ (2)
@@
Looking at Figure 1 we find that the elements utilized in $\mathcal{F}$ are outlined in red. This is continued in all the other Petri net models in this paper to make it easier to find what elements fit where by inspection. For $\mathcal{F}_{1}$ the element $w(t_{N_{j,...,k}},I_{1})=\beta_{1}$. For $\mathcal{F}_{2}$ the element $w(t_{N_{j,...,k}},I_{2})=\beta_{2}$. For $\mathcal{F}_{3}$ no transmission occurs directly into place $I_{3}$, so the element $w(t_{N_{j,...,k}},I_{3})=0$. Thus
@@
We define $w(I_{i},t_{j,...,k})$ as the summation of arc weights of output arcs from infected place $I_{i}$ to transitions $t_{j,...,k}$ for VAPNs and the summation of rate equations of all arcs between infected place $I_{i}$ to transitions $t_{j,...,k}$. In other words, this represents adding up all arcs leaving infected place $I_{i}$. We define $w(t_{I_{i}},I_{j})$ as the arc weights of arcs from transition to infected places where the arcs leading into the transition come from infected place $I_{i}$, and this arc weight is equal to the arc weight of the arc from $t_{I_{i}}$ to $I_{j}$. Lastly, we define $w(t_{N_{m},...,N_{l}},I_{j})$ as the sum of the arc weights of arcs from transitions to infected place $I_{j}$ where the transitions $t_{N_{m},...,N_{l}}$ have no arcs leading into the transitions. We define $V^{-}_{i}(x)$ to be the arc weight of tokens leaving infected place $i$ and $V^{+}_{i}(x)$ to be the arc weight of tokens going into infected place $i$ outside of coming from non-infected places. In terms of what the terms $\mathcal{V}^{-}_{i}(x)$ and $\mathcal{V}^{+}_{i}(x)$ represent, it is the rate of people leaving the infected place $i$ and the rate of people going into infected place $i$ from anything other than coming from non-infected places, respectively. These terms combine to yield, $\mathcal{V}_{i}(x)=\mathcal{V}^{-}_{i}(x)-\mathcal{V}^{+}_{i}(x)$. For Petri nets, this can be expressed through the matrix
$\mathcal{V}_{1,...,T}(x)=\begin{bmatrix}w(I_{1},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{1})&-w(t_{I_{2}},I_{1})&\cdots&-w(t_{I_{T}},I_{1})\\ -w(t_{I_{1}},I_{2})&w(I_{2},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{2})&\cdots&-w(t_{I_{T}},I_{2})\\ \vdots&\vdots&\ddots&\vdots\\ -w(t_{I_{1}},I_{T})&-w(t_{I_{2}},I_{T})&\cdots&w(I_{T},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{T})\end{bmatrix}.$ (3)
Looking at Figure 1 we find that the elements utilized in $\mathcal{V}$ are outlined in green. This is continued in all the other Petri net models in this paper to make it easier to find what elements fit in $\mathcal{V}$ by inspection. For $\mathcal{V}_{1}$ the elements $w(I_{1},t_{j,...,k})=\gamma_{1}$, $w(t_{N_{m},...,N_{l}},I_{1})=0$, $-w(t_{I_{2}},I_{1})=0$, and $-w(t_{I_{3}},I_{1})=0$. For $\mathcal{V}_{2}$ the elements $-w(t_{I_{1}},I_{2})=0$, $w(I_{2},t_{j,...,k})=\gamma_{2}$, $w(t_{N_{m},...,N_{l}},I_{2})=0$, and $--w(t_{I_{3}},I_{2})=0$. For $\mathcal{V}_{3}$ the elements $-w(t_{I_{1}},I_{T})=-\delta_{1}$, $-w(t_{I_{2}},I_{3})=-\delta_{2}$, $w(I_{3},t_{j,...,k})=\varepsilon$, and $w(t_{N_{m},...,N_{l}},I_{3})=0$. Thus,
@@
The disease free equilibrium (DFE) for a system with $L$ places is defined as ($P_{1}^{*},....,P_{L}^{*}$), the state at which the system remains absent of the disease. For the Petri nets the DFE can be found by first setting the token level of all infected places to zero. Then, the arc weights of arcs into and out of a given non-infected place $P_{i}$ are summed, with the the arc weights of arcs going out of $P_{i}$ being multiplied by -1. The resulting expression from these steps is then set to zero. From here, a system of equations is formed that can be solved for $P_{1}^{*},....,P_{L}^{*}$. Additionally, the total population, usually denoted as $N$, is defined as $\mathcal{N}+\varepsilon$, where $\mathcal{N}$ is the summation of tokens in all places, and $\varepsilon$ is the error from rounding in a discrete token level Petri net. In a continuous token level Petri net $\varepsilon=0$. The difference in total population between $N$ and $\mathcal{N}+\varepsilon$ comes into play when doing actual simulations of models not when finding $R_{0}$ using NGMPN method. From here, a system of equations is formed where the DFE token levels labeled for each non-infected place can be solved for.
@@
Then, using the Petri net based $F$ and $V$ values, the same Next-Generation Matrix steps of finding the inverse of $V$ and finding the dominant eigenvalues of $FV^{-1}$ can be done to find the Petri net $R_{0}$. If the values of $F,V$ and the DFE of two systems are equal, then the resulting $R_{0}$ values are equal. However, the way that parameters interact within an ODE model and a PN model are different. We will show the process in detail with various examples with different functionality.
Addressing the assumptions (A1)-(A5) made in van den Driessche and Watmough [6], we interpret them for Petri nets.
@@
    If a place $P_{i}$ contains a non-negative number of tokens $t_{i}\geq 0$, then the number of tokens that can be added to or removed from any place by any transition is non-negative. Specifically, if a transition $t_{j}$ is enabled (based on its input places and corresponding markings), the number of tokens removed from the input places and added to the output places must be non-negative. By the definition of Petri nets, the assumption is inherently preserved across all Petri net structures. This is due to the stipulation that the number of tokens in each place must remain non-negative at all times. Additionally, transitions in a Petri net operate in a manner that ensures the non-negativity of tokens, effectively maintaining this essential characteristic.
@@
    If the marking of a place $P_{i}$ is zero (i.e., $t_{i}=0$), then no transition can consume tokens from place $P_{i}$. Specifically, if a place is empty (i.e., no tokens are present in the corresponding place), no transition can have it as an input. This ensures that no ”removal” of individuals (tokens) from an empty place occurs. Again the definition of Petri net preserves this assumption for all Petri nets as a transition can not fire if an input place has less tokens than the corresponding arc weight. If a place has no tokens (individuals), no transition can remove tokens from that place.
@@
    If the marking of a place $P_{i}$ indicates a disease-free population (i.e., there are no tokens representing infected individuals in this marking), then no transition can fire that would result in tokens (representing infected individuals) being moved to an infected place. If a place is has a disease-free marking, no transitions can place tokens in it that correspond to infected individuals (i.e., no ”infection” transition should have this place as an output if it is currently disease-free).
@@
    If $\mathcal{F}(x)$ is set to 0, then all eigenvalues of $Df(x)$ have negative real parts. In terms of Petri nets, this means that if there are no transitions that change the marking of infected places coming from non-infected places, then the related infected place’s markings will go to zero over time. If a few infected tokens are introduced into the system, the infected place’s markings will still go to zero over time. No new infections means only infections leaving, thus the resulting $Df(x)<0$.
@@ 4 Case Studies
We have selected a range of ODE and corresponding PN example models to show how the Next-Generation Matrix for Petri nets (NGMPN) approach to finding $R_{0}$. Each model either adds additional compartments/places or new dynamics to provide a range of examples for people to draw on if they wish to implement the NGMPN on a new model.
@@ 4.1 SIRS
We start with a simple example of an $SIRS$ system, a small variation on the Kermack-McKendrick model [19]. This shows a straightforward application of the NGMPN process for a foundational epidemiological model. The compartments $S,I,R$ represent susceptible, infected, and recovered, respectively. The parameters of this system $\beta,\delta,$ and $\gamma$ represent the rate of re-susceptibility, rate of infection, and rate of recovery, respectively.
$\frac{dS}{dt}$ $=$ $-\beta SI+\delta R$ (4)
$\frac{dI}{dt}$ $=$ $\beta SI-\gamma I$ (5)
$\frac{dR}{dt}$ $=$ $\gamma I-\delta R$ (6)
The ODE system can be mapped into two Petri net models, one with variable arc weights and minimal arcs yielding Figure 2(a), and another using the methods laid out by Baez and Biamonte [18], yielding Figure2(b).
Figure 2: Two Petri net implementations of the SIRS ODE model given in Equations 4-6. Option (a) uses a discrete-time, variable arc weight PN. Option (b) uses a continuous-time, stochastic Markov Chain PN. The arcs and arc weights used in forming matrix $F_{i}(x)$ are outlined and highlighted in red, and the arcs and arc weights used in forming matrix $V_{i}(x)$ are outlined and highlighted in green.
The first step in finding the $R_{0}$ of the PN systems in Figure 2 is to obtain the Disease Free Equilibrium (DFE) by setting the token levels to zero in the infected places, setting the net arc weight of non-infected places to 0, and solving for the token values in non-infected places $(S^{*},I^{*},R^{*})=(S^{*},0,0)$. Then the transmission matrix starts with finding $\mathcal{F}_{i}(x)$, which can be found by looking at the arc weights of arcs going into the infected compartments as shown in Equation 2. These arcs are outlined in red in Figure 2. Note that $i$ is with respect to the corresponding infected compartment, which we number. For this example, the $I$ compartment is 1. From here we find $w(t_{N_{j,...,k}},I_{1})=\beta SI$ for the VAPN in Figure 2(a) and $w(t_{N_{j,...,k}},I_{1})=2\beta SI-\beta SI$ for the SPN in Figure 2(b). Thus,
$\mathcal{F}_{1}(x)$ $=$ $\begin{bmatrix}\beta SI\end{bmatrix}$
for the Petri net layout as depicted in Figure 2(a), and
@@
for the Petri net layout as depicted in Figure 2(b). From this, we find the rate of change with respect to the infected place $I$, and we evaluate at the DFE to have
$F$ $=$ $\begin{bmatrix}\beta\end{bmatrix}.$
for both layouts. For the transition matrix, we start with $\mathcal{V}$, and we look at arc weights of arcs leaving the infected places as shown in Equation 3. These arcs are outlined in green in Figure 2. The terms $w(I_{1},t_{j,...,k})=\gamma I$ and $w(t_{N_{m},...,N_{l}}=0$ for both the VAPN in Figure 2(a) and the SPN in Figure 2(b). Thus yielding for both layouts
@@
We then find the rate of change with respect to the infected place $I$, and evaluate at the DFE. From this we find the inverse:
$V$ $=$ $\begin{bmatrix}\gamma\end{bmatrix}$
@@
Now that we have the $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and finding the dominant eigenvalue.
$R_{0}$ $=$ $\varphi\begin{bmatrix}FV^{-1}\end{bmatrix}=\frac{\beta}{\gamma}$
This $R_{0}$ value of both layouts of $\frac{\beta}{\gamma}$ is analytically equivalent to the ODE $R_{0}$ value.

### v3

@@ 3.2 Constructing the Next-Generation Matrix for Petri Nets
Figure 1: This VAPN is used as a guide to help visualize some of the concepts described in the Next-Generation Matrix for Petri Nets formulation. The legend on the right describes the elements within the PN. The places of this VAPN can be labeled $(P_{1},P_{2},P_{3},P_{4},P_{5})$ or equivalently $(N_{1},I_{1},I_{2},I_{3},N_{2})$, where the second label gives more description on the infectious state of the place and is used in the diagram. Red outline correspond to use in the corresponding $\mathcal{F}$ matrix and green outline correspond to use in the corresponding $\mathcal{V}$ matrix.
@@
In a Petri Net, places are the nodes that hold tokens, representing subpopulations or the state of the system. Then, for the purpose of our framework, it is helpful to think of places in Petri Nets the same way that you might think about compartments in ODEs. Note that in systems biology a “compartment” in a Petri Net is a way to group related places and transitions within a larger Petri Net Heiner et al. (2009). Then, we assume we have a Petri Net model for a disease that fits the definition of VAPN or SPN outlined in Section 2. Assume that each transition can fire once per time step. This assumption does not lead to a loss of generality because the framework’s core calculation relies on the rate of change in token flow, which applies to both the functional arc weights in deterministic VAPNs and the continuous-time firing rates in SPNs. The “time step” serves as a conceptual unit for a state update rather than a rigid constraint on the model’s dynamics. Define the total number of infected places in the Petri Net model as $Q$. Thus, the infected places are defined as $I_{1},I_{2},\ldots,I_{Q}$. As a note, we define all places as $P$ and they are labeled $P_{1},\ldots,P_{L}$, where $L$ is the total number of places and $Q\leq L$. We define $t_{N}$ as a transition with an input arc coming from a non-infected place and an output arc going to an infected place. With $w(t_{N_{j,\ldots,k}},I_{i})$ being the sum of all arcs weights $j,\ldots,k$ from arcs going from non-infected places going to an infected place $I_{i}$ for VAPN or the net sum of all rate equations $j,\ldots,k$ for arcs from transitions with input arcs from non-infected places and output arcs going to infected places $I_{i}$ for SPN. Let $x=(x_{1},…,x_{n})^{t}$, with each $x_{i}\geq 0$, be the number of tokens in each place, which is the Petri Net equivalent of population in each compartment in an ODE. Then we define $\mathcal{F}_{i}(x)$ to be the rate at which tokens move into infected places. For the variable arc weight PN, this rate refers to the arc weights.
The following provides the clear formal algorithm for distinguishing new infection arcs ($F$) from other state change arcs ($V$). First, let the set of all places be partitioned into infected places, $P_{I}$, and non-infected places, $P_{N}$. We then define an infection transition $t\in T$ as any transition that has at least one input arc from a place in $P_{N}$ and at least one output arc to a place in $P_{I}$. Let the set of all such infection transitions be $T_{inf}$. From this formal definition, we construct the matrices $F$ and $V$:
  • For $F$: For any infected place $p_{i}\in P_{I}$, we define $\mathcal{F}_{i}(x)$ as the rate of all new infections entering $p_{i}$, which is the net inflow from all infection transitions $t_{j}\in T_{inf}$. The matrix $\mathcal{F}(x)$ is composed of these $\mathcal{F}_{i}(x)$ terms.
  • For $V$: Conversely, $\mathcal{V}_{i}(x)$ represents the net rate of all other transfers affecting place $p_{i}$. This is calculated as the rate of all outflows from $p_{i}$ (to any place) minus the rate of all inflows to $p_{i}$ from other infected places $p_{k}\in P_{I}$.
The matrices in Equations (2) and (3) provide the operational method for calculating these conceptual rates. Then $\mathcal{F}_{i}(x)$ is represented in matrix form by
$\mathcal{F}_{1,\ldots,T}(x)$ $=\begin{bmatrix}w(t_{N_{j,\ldots,k}},I_{1})&w(t_{N_{j,\ldots,k}},I_{1})&\cdots&w(t_{N_{j,\ldots,k}},I_{1})\\ w(t_{N_{j,\ldots,k}},I_{2})&w(t_{N_{j,\ldots,k}},I_{2})&\cdots&w(t_{N_{j,\ldots,k}},I_{2})\\ \vdots&\vdots&\ddots&\vdots\\ w(t_{N_{j,\ldots,k}},I_{T})&w(t_{N_{j,\ldots,k}},I_{T})&\cdots&w(t_{N_{j,\ldots,k}},I_{T}))\end{bmatrix}$. (2)
@@
Examining Figure 1, we observe that the parameters of the PN that correspond to the $\mathcal{F}$ matrix entries are outlined in red. This is continued in all the other Petri Net models in this paper to facilitate easier identification of elements by inspection. For $\mathcal{F}_{1}$ the element $w(t_{N_{j,\ldots,k}},I_{1})=\beta_{1}$. For $\mathcal{F}_{2}$ the element $w(t_{N_{j,\ldots,k}},I_{2})=\beta_{2}$. For $\mathcal{F}_{3}$ no transmission occurs directly into place $I_{3}$, so the element $w(t_{N_{j,\ldots,k}},I_{3})=0$. Thus
@@
We define $w(I_{i},t_{j,\ldots,k})$ as the summation of arc weights of output arcs from infected place $I_{i}$ to transitions $t_{j,\ldots,k}$ for VAPNs and the summation of rate equations of all arcs between infected place $I_{i}$ to transitions $t_{j,\ldots,k}$. In other words, this represents adding up all arcs leaving the infected place $I_{i}$. We define $w(t_{I_{i}},I_{j})$ as the arc weights of arcs from transition to infected places where the arcs leading into the transition come from infected place $I_{i}$, and this arc weight is equal to the arc weight of the arc from $t_{I_{i}}$ to $I_{j}$. Lastly, we define $w(t_{N_{m},\ldots,N_{l}},I_{j})$ as the sum of the arc weights of arcs from transitions $t_{N_{m},\ldots,N_{l}}$ to infected place $I_{j}$, where these transitions have no arcs leading into them. We define $V^{-}_{i}(x)$ to be the arc weight of tokens leaving infected place $i$ and $V^{+}_{i}(x)$ to be the arc weight of tokens going into infected place $i$ outside of coming from non-infected places. In terms of what the terms $\mathcal{V}^{-}_{i}(x)$ and $\mathcal{V}^{+}_{i}(x)$ represent, it is the rate of people leaving the infected place $i$ and the rate of people going into infected place $i$ from anything other than coming from non-infected places, respectively. These terms combine to yield, $\mathcal{V}_{i}(x)=\mathcal{V}^{-}_{i}(x)-\mathcal{V}^{+}_{i}(x)$. For Petri Nets, this can be expressed through the matrix
$\mathcal{V}_{1,\ldots,T}(x)=\begin{bmatrix}w(I_{1},t_{j,\ldots,k})-w(t_{N_{m},\ldots,N_{l}},I_{1})&-w(t_{I_{2}},I_{1})&\cdots&-w(t_{I_{T}},I_{1})\\ -w(t_{I_{1}},I_{2})&w(I_{2},t_{j,\ldots,k})-w(t_{N_{m},\ldots,N_{l}},I_{2})&\cdots&-w(t_{I_{T}},I_{2})\\ \vdots&\vdots&\ddots&\vdots\\ -w(t_{I_{1}},I_{T})&-w(t_{I_{2}},I_{T})&\cdots&w(I_{T},t_{j,\ldots,k})-w(t_{N_{m},\ldots,N_{l}},I_{T})\end{bmatrix}$. (3)
[…]
@@
The disease free equilibrium (DFE) for a system with $L$ places is defined as ($P_{1}^{*},\ldots,P_{L}^{*}$), the state at which the system remains free of the disease. For the Petri Nets, the DFE can be found by first setting the token level of all infected places to zero. Then, the arc weights of arcs into and out of a given non-infected place $P_{i}$ are summed, with the arc weights of arcs going out of $P_{i}$ being multiplied by -1. The resulting expression from these steps is then set to zero. From here, a system of equations is formed that can be solved for $P_{1}^{*},\ldots,P_{L}^{*}$. Additionally, the total population, usually denoted as $N$, is defined as $\mathcal{N}+\varepsilon$, where $\mathcal{N}$ is the summation of tokens in all places, and $\varepsilon$ is the error from rounding in a discrete token level Petri Net. In a continuous token level Petri Net, $\varepsilon=0$. The difference in total population between $N$ and $\mathcal{N}+\varepsilon$ becomes relevant when performing actual simulations of models, rather than when finding $R_{0}$ using the NGMPN method. From here, a system of equations is formed where the DFE token levels labeled for each non-infected place can be solved for.
@@
Then, using the Petri Net-based $F$ and $V$ values, the same Next-Generation Matrix steps of finding the inverse of $V$ and finding the dominant eigenvalues of $FV^{-1}$ can be done to find the Petri Net $R_{0}$. If the values of $F,V$, and the DFE of two systems are equal, then the resulting $R_{0}$ values are equal. However, the way parameters interact within an ODE model and a PN model differ. We will demonstrate the process in detail, providing various examples with different functionalities.
Addressing the assumptions (A1)–(A5) made in van den Driessche and Watmough Van den Driessche and Watmough (2002), we interpret them for Petri Nets.
@@
    If a place $P_{i}$ contains a non-negative number of tokens $t_{i}\geq 0$, then the number of tokens that can be added to or removed from any place by any transition is non-negative. Specifically, if a transition $t_{j}$ is enabled (based on its input places and corresponding markings), the number of tokens removed from the input places and added to the output places must be non-negative. By definition of Petri Nets, the assumption is inherently preserved across all Petri Net structures. This is due to the stipulation that the number of tokens in each place must remain non-negative at all times. Additionally, transitions in a Petri Net operate in a manner that ensures the non-negativity of tokens, effectively maintaining this essential characteristic.
@@
    If the marking of a place $P_{i}$ is zero (i.e., $t_{i}=0$), then no transition can consume tokens from place $P_{i}$. Specifically, if a place is empty (i.e., no tokens are present in the corresponding place), no transition can have it as an input. This ensures that no “removal” of individuals (tokens) from an empty place occurs. Again, the definition of Petri Net preserves this assumption for all Petri Nets, as a transition can not fire if an input place has less tokens than the corresponding arc weight. If a place has no tokens (individuals), no transition can remove tokens from that place.
@@
    If the marking of a place $P_{i}$ indicates a disease-free population (i.e., there are no tokens representing infected individuals in this marking), then no transition can fire that would result in tokens (representing infected individuals) being moved to an infected place. If a place has a disease-free marking, no transitions can place tokens in it that correspond to infected individuals (i.e., no “infection” transition should have this place as an output if it is currently disease-free).
@@
    If $F(x)$ is set to 0, then all eigenvalues of the Jacobian of the infected subsystem at the DFE have negative real parts. In terms of the Petri Nets, this means that when no new infections occur, that means all transitions transferring tokens from non-infected to infected places are disabled, the marking of every infected place will asymptotically decay to zero. Formally, let the state vector be partitioned as $x=(x_{I},x_{S})$, where $x_{I}$ are the infected places and $x_{S}$ are the non-infected places. Define the infected subsystem as $\dot{x}_{I}=F_{I}(x_{I},x_{S})-V_{I}(x_{I},x_{S})$, where $F_{I}$ collects all token inflows into infected places from non-infected places and $V_{I}$ collects all other token flows (recoveries, deaths, or transfers among infected places). Then, at the DFE $x_{0}$, with $F_{I}(x_{0})=0$, the Jacobian matrix
    $A=D_{x_{I}}(F_{I}-V_{I})\big|_{x=x_{0}}=-D_{x_{I}}V_{I}(x_{0})$
    must be Hurwitz, which means all eigenvalues have strictly negative real parts. Note that $D$ is the derivative operator. Equivalently, $D_{x_{I}}V_{I}(x_{0})$ should be a nonsingular $M$-matrix, ensuring that small perturbations in the infected places decay over time. In terms of a biological application, this assumption guarantees that in the absence of new infections, any initial infection will die out, preserving the disease-free equilibrium.
@@ 4 Case Studies
We have selected a range of ODE and corresponding PN example models to demonstrate how the Next-Generation Matrix for Petri Nets approach to determining $R_{0}$ works. Each model either adds additional compartments/places or introduces new dynamics, providing a range of examples for people to draw on if they wish to implement the NGMPN in a new model.
@@ 4.1 SIRS
We start with a simple example of an $SIRS$ system, a slight variation in the Kermack–McKendrick model Kermack and McKendrick (1927). This shows a straightforward application of the NGMPN process for a foundational epidemiological model. The compartments $S$, II, and $R$ represent susceptible, infected, and recovered, respectively. The parameters of this system $\beta,\delta,$ and $\gamma$ represent the rate of re-susceptibility, rate of infection, and rate of recovery, respectively. The SIRS system is described by the following system of equations:
$\frac{dS}{dt}$ $=-\beta SI+\delta R,$ (4)
$\frac{dI}{dt}$ $=\beta SI-\gamma I,$ (5)
$\frac{dR}{dt}$ $=\gamma I-\delta R.$ (6)
The ODE system can be mapped into two Petri Net models: one with variable arc weights and minimal arcs, yielding Figure 2a, and another using the methods outlined by Baez and Biamonte Baez and Biamonte (2012), yielding Figure 2b.
Figure 2: Two Petri Net implementations of the SIRS ODE model given in Equations (4)–(6). Option (a) uses a discrete-time, variable arc weight PN. Option (b) uses a continuous-time, stochastic Markov Chain PN. The arcs and arc weights used in forming matrix $F_{i}(x)$ are outlined and highlighted in red, and the arcs and arc weights used in forming matrix $V_{i}(x)$ are outlined and highlighted in green.
The first step in finding the $R_{0}$ of the PN systems in Figure 2 is to obtain the Disease Free Equilibrium (DFE) by setting the token levels to zero in the infected places, setting the net arc weight of non-infected places to 0, and solving for the token values in non-infected places $(S^{*},I^{*},R^{*})=(S^{*},0,0)$, where $S^{*}=N$. Where $N=S+I+R$ or the sum of the entire population. Then the transmission matrix starts with finding $\mathcal{F}_{i}(x)$, which can be found by looking at the arc weights of arcs going into the infected compartments as shown in Equation (2). These arcs are outlined in red in Figure 2. Note that $i$ refers to the corresponding infected compartment, which we have numbered. For this example, the $I$ compartment is 1. From here we find $w(t_{N_{j,\ldots,k}},I_{1})=\beta SI$ for the VAPN in Figure 2a and $w(t_{N_{j,\ldots,k}},I_{1})=2\beta SI-\beta SI$ for the SPN in Figure 2b. Thus,
$\mathcal{F}_{1}(x)$ $=\begin{bmatrix}\beta SI\end{bmatrix}$,
for the Petri Net layout as depicted in Figure 2a, and
@@
for the Petri Net layout as depicted in Figure 2b. From this, we find the rate of change with respect to the infected place $I$, and we evaluate it at the DFE to have
$F$ $=\begin{bmatrix}\beta N\end{bmatrix}$.
for both layouts. For the transition matrix, we start with $\mathcal{V}$, and we look at arc weights of arcs leaving the infected places as shown in Equation (3). These arcs are outlined in green in Figure 2. The terms $w(I_{1},t_{j,\ldots,k})=\gamma I$ and $w(t_{N_{m},\ldots,N_{l}}=0$ for both the VAPN in Figure 2a and the SPN in Figure 2b. Thus yielding for both layouts
@@
We then find the rate of change with respect to the infected place $I$ and evaluate at the DFE. From this, we find the inverse:
$V$ $=\begin{bmatrix}\gamma\end{bmatrix}$,
@@
Now that we have the $F$ and $V$ matrices, we can find $R_{0}$ by multiplying them and then determining the dominant eigenvalue we get,
$R_{0}$ $=\varphi\begin{bmatrix}FV^{-1}\end{bmatrix}=\frac{\beta N}{\gamma}.$
This $R_{0}$ value of both layouts of $\frac{\beta N}{\gamma}$ is analytically equivalent to the ODE $R_{0}$ value.